solve the inequality,state the solution set using interval notation and graph the solution set.
4x^2-8x greater than or equal to 0

Answer by stop12345now(6)

4x^2-8x > 0 (add equals after solution)
4x(x-2) > 0 (factor out 4x)
create [plus and minus] chart
4x>0
//////////////////////////////
________________________________________
012345678
x-2>0
zero is at2
//////////////////////////// _________________________
2345678
u may have to fill in negative numbers for chart because computer messes it up.
also reposition the two so it matches the two in the First chart.
draw lines through the zeros (0 for top, 2 for bottom)
mark the signs of each zone
multiply top and bottom zones sep. by lines.
solution : whenever signs are the same
solution
2 -infinity [2,infinity)
(-infinity, 0)
the graph is a parabola
vertex -b/2a=-(-8/4*2)=1 (x coordinate)
y=4(1)^2-8(1)=4-8=-4
the graph is a parabola with vertex (1,-4)pointing up like a "U"